# SpecServe: Efficient and SLO-Aware Large Language Model Serving with Adaptive Speculative Decoding

### v1

@@ 2.4 Key Insight and Challenges
Key Insight: Fine-grained dynamic control. These limitations discussed in §2.3 reveal a core contradiction: the potential benefits of speculative decoding are constrained by its rigid execution framework. We propose a fundamental solution which transforms the speculative length from a preset parameter into a dynamically controlled variable. Specifically, the system needs to dynamically adjust the speculative length for each decoding step in real-time, based on workload characteristics, system configurations, and SLO constraints. This mechanism enables speculative decoding to both mitigate the risk of resource overload (e.g., by shortening speculative lengths under high-load conditions) and maximize computational resource utilization (e.g., by extending speculative lengths under low-load conditions), thereby overcoming the performance boundaries of static strategies.
However, achieving fine-grained dynamic control of speculative length faces three major challenges.
@@
Challenge #2: How to timely and flexibly control the speculative length? The root of this challenge lies in two main factors: (1) Lag in efficiency estimation: The estimation of decoding efficiency depends on the actual execution results of the draft model. This dependency prevents the system from preemptively evaluating the validity of a given speculative length before generating the draft, leading to delay in controlling. (2) Rigid system architecture constraints: All requests within the same batch have the same number of draft model executions, which results in the default speculative length being the same for all requests, thereby neglecting fine-grained control.
@@
Challenge #3: How to balance the efficiency of speculative decoding and SLO attainment? Due to the probabilistic fluctuations in the acceleration effects of speculative decoding, there is a conflict between maximizing the efficiency of speculative decoding and meeting SLO requirements (§2.3). Therefore, we need to incorporate SLO constraints and the risk of TPOT exceeding expectations into the efficiency model, allowing the system to balance acceleration gains and SLO violation risks in real-time when controlling speculative length.
@@
Figure 6: An overview of SpecServe’s modules: adaptive drafter, confidence prior verifier, and SLO-aware efficiency estimator.
3 SpecServe’s Overview
@@
We proposed SpecServe, an efficient LLM inference system with adaptive speculative decoding. To accurately characterize the dynamic relationship between efficiency and speculative length (Challenge #1), we establish an effective and general-purpose efficiency model for speculative decoding (§4). Based on this model, we implement three components to achieve fine-grained speculative decoding, as shown in Fig. 6.
@@
Adaptive Drafter (§5.1). To address the lag in efficiency estimation (Challenge #2-1), SpecServe introduces an adaptive drafter, integrating efficiency estimation into the drafting phase to achieve proactive, step-level speculative length control. As illustrated in Fig. 6, before each execution of the draft model, the speculative controller uses historical confidence scores to estimate the potential efficiency gain of the current speculative length in real-time. After executing the draft model, the generated confidence scores are recorded and used to adjust the efficiency of the current length to ensure accurate speculative length control.
@@
Confidence Prior Verifier (§5.2). To address the rigid system architecture constraints (Challenge #2-2), SpecServe employs a confidence prior verifier to prioritize the verification of tokens with high acceptance rates, thereby decoupling the drafting phase from the verification phase and enabling fine-grained request-level speculative length control. As shown in Fig. 6, before being input into the target model, the draft token set undergoes draft token elimination, where tokens are progressively removed in ascending order of acceptance rate until the draft token set with the highest efficiency is selected.
@@
SLO-Aware Efficiency Estimator (§5.3). The key operations of SpecServe are based on the evaluation of the efficiency of speculative decoding. Similar to vLLM [17] and Orca [38], we use throughput to measure decoding efficiency and propose an efficiency estimator. This estimator employs the efficiency model of speculative decoding to adapt to dynamic request patterns and diverse system configurations. To achieve SLO awareness (Challenge #3), the efficiency estimator incorporates SLO constraints to support rapid trade-offs between the efficiency of speculative decoding and SLO compliance.
4 Modeling the Efficiency of Speculative Decoding
@@
To evaluate the efficiency of speculative decoding, we adopted the same throughput metric as used in existing LLM inference systems [17, 38], which is defined as $Throughput=NAT/T_{sd}$. We developed two models: one for speculative decoding time ($T_{sd}$) and another for the number of accepted tokens ($NAT$). Based on these models, we investigated the relationship between the efficiency of speculative decoding and the speculative length, finding that the efficiency of speculative decoding initially increases and then decreases, or it monotonically decreases, with an increase in speculative length. This discovery provides a theoretical foundation for implementing efficient speculative length control.
4.1 The Time for Speculative Decoding
@@
The time for speculative decoding $T_{sd}$ includes the execution time of the draft model ($T_{d}$) and the execution time of the target model ($T_{v}$), which can be expressed as:
$T_{sd}=T_{d}+T_{v}$ (1)
Modeling the forward propagation time. The execution of both the draft and the target models fundamentally involves multiple instances of forward propagation. Similar to previous work [22], we simplify the time for forward propagation as:
$T_{fwd}(N_{c},N_{b})=\alpha*N_{c}+\gamma*N_{b}+\delta$ (2)
, where $N_{c}$ and $N_{b}$ respectively represent the number of context tokens and batch tokens in the batch, and the performance coefficients $(\alpha,\gamma,\delta)$ depend on the computing platform, parallelism, and the model.
Modeling the execution time of draft and target models. The execution time of the draft model is the sum of the times for multiple forward passes, which can be expressed as:
$T_{d}$ $=\sum_{i=1}^{SL}\big{(}\alpha_{d}*N_{dc}(i)+\gamma_{d}*N_{db}(i)+\delta_{d}\big{)}$ (3)
$=\alpha_{d}*\sum_{i=1}^{SL}N_{dc}(i)+\gamma_{d}*\sum_{i=1}^{SL}N_{db}(i)+\delta_{d}*SL$ (4)
, where $(\alpha_{d},\gamma_{d},\delta_{d})$ is the performance coefficients of the draft model on the given computing platform, and $N_{dc}(i),N_{db}(i)$ are the number of context tokens and batch tokens in the batch during the $i^{th}$ forward propagation of the draft model.
Similarly, the execution time of the target model can be represented as the time for one forward propagation to process all speculative inputs, as shown in Eq. 5. We can relate the number of context tokens $N_{vc}$ and batch tokens $N_{vb}$ in each batch of the target model’s forward propagation to $N_{dc}$ and $N_{db}$, as shown in Eq. 6.
$T_{v}$ $=\alpha_{v}*N_{vc}+\gamma_{v}*N_{vb}+\delta_{v}$ (5)
$=\alpha_{v}*\sum_{i=1}^{SL+1}N_{dc}(i)+\gamma_{v}*\sum_{i=1}^{SL+1}N_{db}(i)+\delta_{v}$ (6)
@@
, where $(\alpha_{v},\gamma_{v},\delta_{v})$ is the performance coefficients of the target model on the given computing platform.
@@
Figure 7: Average acceptance rates for different confidence score intervals of the draft model. The red dashed line connects (0,0) and (1,1) to aid in visual assessment.
4.2 The Number of Accepted Tokens
Given the dynamic system loads and diverse request categories, it is impractical to estimate the number of accepted tokens for any given request using overly complex conditions. Therefore, we propose an acceptance token number model that does not rely on any conditions:
$NAT=\sum_{i=1}^{BS}(\sum_{j=1}^{SL}AR_{i,j}+1)=\sum_{i=1}^{BS}(\sum_{j=0}^{SL}AR_{i,j})$ (7)
, where $AR_{i,j}$ denotes the probability that the $j^{th}$ draft token of the $i^{th}$ request is accepted in the current step. Since every request has a bonus token that is inevitably accepted, the sum of the acceptance probabilities must also include this bonus token. In fact, autoregressive decoding can be considered as speculative decoding with a speculative length $SL=0$, accepting only one bonus token. To unify autoregressive decoding with speculative decoding, we define $AR_{i,0}=1$.
Figure 8: The relationship between estimated and actual number of accepted token with speculative length.
Modeling acceptance rate. The experimental results shown in Fig. 7 demonstrate that the acceptance rate of draft tokens under different combinations of draft and target models is positively correlated with their confidence scores in the draft model. Similar to previous works [8, 20], this finding suggests that we can model acceptance rates using confidence levels.
Considering the sequential dependency among tokens, we represent the probability that the $j^{th}$ draft token of the $i^{th}$ request is accepted in the current speculative decoding step as the probability that all previous $j$ draft tokens are accepted simultaneously [37], which is:
$AR_{i,j}=\prod_{k=0}^{j}p_{i,k}\approx\prod_{k=0}^{j}c_{i,k}$ (8)
@@ 4.2 The Number of Accepted Tokens
Here, $p_{i,k}$ is the conditional probability of the $k^{th}$ draft token of the $i^{th}$ request is accepted given that all previous draft tokens have been accepted, and $c_{i,k}$ is its confidence score. Notably, we define $p_{i,0}=c_{i,0}=1$ to unify autoregressive decoding and speculative decoding.
The effectiveness of acceptance rate. Based on Eq. 8, the acceptance rate $AR$ exhibits two mathematical properties:
  1. 1.
    $AR\in[0,1]$, consistent with the nature of probability.
  2. 2.
    $AR_{i,j+1}\leq AR_{i,j}$, because $AR_{i,j+1}=AR_{i,j}*p_{i,j+1}$, and $p_{i,j+1}\leq 1$
At the same time, if the acceptance or rejection of draft tokens during the verification phase corresponds to an acceptance rate of 1 or 0, respectively, then the actual acceptance situation also conforms to the properties modeled by the acceptance rate, demonstrating the effectiveness of acceptance rate modeling in depicting actual acceptance situations.
$NAT=BS*\sum_{j=0}^{SL}\overline{AR_{j}}=BS*\overline{NAT}$ (9)
@@
Further, as shown in Eq. 9, the mathematical properties of the average number of accepted tokens per request $\overline{NAT}$ can be derived as follows:
  1. 1.
    When $SL=0$, $\overline{NAT}=1$.
  2. 2.
    When $SL$ increases, $\Delta\overline{NAT}=\overline{AR}\in[0,1]$
  3. 3.
@@
    When $SL$ increases, $\Delta\overline{NAT}$ is non-increasing.
@@
4.3 The Relationship between Efficiency and Speculative Length
To achieve optimal efficiency through fine-grained control of speculative length, it is first necessary to understand the relationship between efficiency and speculative length, that is, how throughput varies with speculative length.
@@ 4.3 The Relationship between Efficiency and Speculative Length
The relationship between execution time and speculative length. Given the average number of context tokens per request $\overline{N_{c}}$ and the batch size $BS$, we can derive $N_{dc}(i)$ and $N_{db}(i)$. Then, the speculative decoding time is given by:
$\begin{split}T_{sd}&=BS\big{(}\frac{SL^{2}(\alpha_{d}+\alpha_{v})}{2}\\ &\quad+SL(\alpha_{d}\overline{N_{c}}+\alpha_{v}\overline{N_{c}}-\frac{\alpha_{d}}{2}+\frac{\alpha_{v}}{2}+\gamma_{d}+\gamma_{v}+\frac{\delta_{d}}{BS})\\ &\quad+\alpha_{v}\overline{N_{c}}+\gamma_{v}+\frac{\delta_{v}}{BS}\big{)}\end{split}$ (10)
It is evident that the speculative decoding time is a quadratic function of the speculative length $SL$.
The relationship between throughput and speculative length. Throughput can be expressed as a function of speculative length, denoted as $f(s)=g(s)/h(s)$. Here, $g(s)$ represents the average number of tokens received per request ($\overline{NAT}$), and $h(s)$ represents the average speculative decoding time per request ($T_{sd}/{BS}$), where $s$ is the continuous speculative length. Herein, $g(s)$ and $h(s)$ possess the following properties:
  1. 1.
    $g(0)=1$, $0\leq g^{\prime}(s)\leq 1$, $g^{\prime\prime}(s)\leq 0$
  2. 2.
    $h(s)=as^{2}+bs+c>0$, $h^{\prime}(s)=2as+b>0$, $h^{\prime\prime}(s)=c>0$
The first derivative $f^{\prime}(s)$ and the second derivative $f^{\prime\prime}(s)$ of $f(s)$ can be represented as follows:
$f^{\prime}(s)=\frac{g^{\prime}(s)h(s)-g(s)h^{\prime}(s)}{\big{(}h(s)\big{)}^{2}}$ (11)
$\begin{split}f^{\prime\prime}(s)=&\frac{1}{\big{(}h(s)\big{)}^{4}}\Big{(}\big{(}g^{\prime\prime}(s)h(s)-g(s)h^{\prime\prime}(s)\big{)}\big{(}h(s)\big{)}^{2}\\ &-2h(s)h^{\prime}(s)\big{(}g^{\prime}(s)h(s)-g(s)h^{\prime}(s)\big{)}\Big{)}\end{split}$ (12)
It is evident that when $f^{\prime}(s)>0$, $f^{\prime\prime}(s)<0$. This characteristic indicates that as $s$ increases, $f^{\prime}(s)$ will eventually change within the negative range. Moreover, when $f^{\prime}(0)>0$, that is, when $g^{\prime}(0)c-b>0$, $f(s)$ initially increases and then decreases; otherwise, $f(s)$ monotonically decreases.
Furthermore, the variation of $f(s)$ is determined by:
@@
$\begin{split}g^{\prime}(0)c-b=&2\alpha_{v}\overline{N_{c}}g^{\prime}(0)+2\gamma_{v}g^{\prime}(0)+\frac{2\delta_{v}}{BS}g^{\prime}(0)-2\alpha_{v}\overline{N_{c}}\\ &-\alpha_{v}-2\gamma_{v}-2\alpha_{d}\overline{N_{c}}+\alpha_{d}-2\gamma_{d}-\frac{2\delta_{d}}{BS}\end{split}$ (13)
, implying that throughput tends to first increase and then decrease over speculative length more readily under conditions including higher acceptance rates with greater $g^{\prime}(0)$, smaller batch sizes with lower $BS$, shorter contexts with fewer $\overline{N_{c}}$, larger target models with higher values of $(\alpha_{v},\gamma_{v},\delta_{v})$, and smaller draft models with lower values of $(\alpha_{d},\gamma_{d},\delta_{d})$.
In summary, we prove that there exists a unique optimal speculative length that can maximize the efficiency of speculative decoding. Therefore, we do not need to preset a range and iterate over all possible lengths to calculate the efficiency of speculative decoding; we only need to find this unique peak value.
5 Adaptive Speculative Decoding

### v2

@@ 2.4. Key Insight and Challenges
Key Insight: Fine-grained dynamic control. These limitations discussed in §2.3 reveal a core contradiction: the potential benefits of speculative decoding are constrained by its rigid execution framework. We propose a solution which transforms the speculative length from a preset parameter into a dynamically controlled variable. Specifically, the system needs to dynamically adjust the speculative length for each decoding step in real-time, based on workload characteristics, system configurations, and SLO constraints. This resource scheduling mechanism enables speculative decoding to both mitigate the risk of resource overload (e.g., by shortening speculative lengths under high-load conditions) and maximize computational resource utilization (e.g., by extending speculative lengths under low-load conditions), thereby overcoming the performance boundaries of static strategies. However, achieving fine-grained dynamic control of speculative length faces three major challenges.
@@
Figure 5. Relationship between SLO attainment and end-to-end latency speedup under different speculative lengths. The red dashed line represents the speedup value of 1 and the SLO attainment value of 0.9.
@@
Challenge #2: How to timely and flexibly control the speculative length? The root of this challenge lies in two factors: (1) Lag in efficiency estimation: The estimation of decoding efficiency depends on the execution results of the draft model. This dependency prevents the system from preemptively evaluating the validity of a given speculative length before generating the draft, leading to delay in controlling. (2) Rigid system architecture constraints: All requests within the same batch have the same number of draft model executions, which results in the speculative length being the same for all requests, thereby neglecting fine-grained control.
@@
Challenge #3: How to balance the efficiency of speculative decoding and SLO attainment? Due to probabilistic fluctuations in the acceleration effects of speculative decoding, there is a conflict between maximizing the efficiency of speculative decoding and meeting SLO requirements (§2.3). Therefore, we need to incorporate SLO constraints and the risk of TPOT exceeding expectations into the efficiency model, to balance acceleration gains and SLO violation risks in real-time when controlling speculative length.
@@
Figure 6. An overview of AdaSpec’s modules: adaptive drafter, confidence prior verifier, and SLO-aware efficiency estimator.
3. AdaSpec’s Overview
@@
We proposed AdaSpec, an efficient LLM inference system with adaptive speculative decoding. To accurately characterize the dynamic relationship between efficiency and speculative length (Challenge #1), we establish an effective and general-purpose efficiency model for speculative decoding (§4). Building upon this foundation, AdaSpec adopts a modular design comprising three key components: adaptive drafter, confidence prior verifier, and SLO-aware efficiency estimator. The design principles of these components and their collaborative operational mechanism are illustrated in Fig. 6.
@@
Adaptive Drafter (§5.1). To achieve timely control over speculative length (Challenge #2-1), AdaSpec introduces an adaptive drafter, integrating efficiency estimation into the drafting phase to achieve real-time, step-level speculative length control. As illustrated in Fig. 6, the adaptive drafter consists of a speculative controller and a draft model governed by it. Before each execution of the draft model, the speculative controller uses historical confidence scores to estimate the potential efficiency gain of the current speculative length in real-time. The draft model executes only if a positive gain is predicted; otherwise, the execution terminates timely. After executing the draft model, the generated confidence scores are recorded as historical data and used to calibrate the actual efficiency gain of the current length, ensuring precise control over speculative length.
@@
Confidence Prior Verifier (§5.2). To enable flexible control over speculative length (Challenge #2-2), AdaSpec employs a confidence prior verifier to prioritize the verification of tokens with high acceptance rates. This approach decouples the dependency between the drafting and verification phases, breaking the rigid system architecture constraints and achieving fine-grained, request-level speculative length control. As shown in Fig. 6, the confidence prior verifier consists of a draft token elimination and a target model. Before being input into the target model, the draft token set generated by the adaptive drafter undergoes draft token elimination, where tokens are removed in ascending order of acceptance rate until the draft token set with the highest efficiency is selected.
@@
SLO-Aware Efficiency Estimator (§5.3). The key operations of AdaSpec (speculative controller and draft token elimination) are based on the evaluation of the efficiency of speculative decoding. Similar to vLLM (Kwon et al., 2023) and Orca (Yu et al., 2022), we use throughput to measure decoding efficiency and propose an efficiency estimator. This estimator employs the efficiency model of speculative decoding to adapt to dynamic request patterns and diverse system configurations. To achieve SLO awareness (Challenge #3), the efficiency estimator incorporates SLO constraints to support rapid trade-offs between the efficiency of speculative decoding and SLO compliance.
4. Modeling the Efficiency of Speculative Decoding
@@
To evaluate the efficiency of speculative decoding, we adopted the same throughput metric as used in LLM inference systems (Kwon et al., 2023; Yu et al., 2022), which is defined as $Throughput=NAT/T_{sd}$. By analyzing how the number of accepted tokens ($NAT$) and the time of speculative decoding ($T_{sd}$) vary with speculative length, we show that throughput either peaks at an intermediate length or decreases monotonically, providing a theoretical basis for efficient length tuning.
4.1. The Time for Speculative Decoding
@@
The time for speculative decoding $T_{sd}$ includes the execution time of the draft model ($T_{d}$) and the target model ($T_{v}$), expressed as:
(1) $T_{sd}=T_{d}+T_{v}$
Modeling the forward propagation time. Execution time in transformer models is dominated by forward passes. Following  (Liu et al., 2024b), we model the forward propagation cost for a batch as a linear function of the number of context tokens $N_{c}$ and batch tokens $N_{b}$.
(2) $T_{fwd}(N_{c},N_{b})=\alpha*N_{c}+\gamma*N_{b}+\delta$
@@
, where the performance coefficients $(\alpha,\gamma,\delta)$ depend on the computing platform, parallelism, and the model.
Modeling the execution time of draft and target models. The draft model performs $SL$ forward passes and the target model performs one forward pass over all $SL+1$ positions. Under Eq.  (2), the execution time of draft and target models can be expressed as:
(3) $T_{d}$ $=\alpha_{d}*\sum_{i=1}^{SL}N_{c}(i)+\gamma_{d}*\sum_{i=1}^{SL}N_{b}(i)+SL*\delta_{d}$
(4) $T_{v}$ $=\alpha_{v}*\sum_{i=1}^{SL+1}N_{c}(i)+\gamma_{v}*\sum_{i=1}^{SL+1}N_{b}(i)+\delta_{v}$
Here, $N_{c}(i)$ and $N_{b}(i)$ denote the number of context tokens and batch tokens in the $i^{th}$ forward pass of the draft model, and $(\alpha_{d},\gamma_{d},\delta_{d})$ and $(\alpha_{v},\gamma_{v},\delta_{v})$ are the performance coefficients for the draft and target models, respectively.
Modeling the relationship with speculative length. To understand how speculative length $SL$ influences overall latency, we express $T_{sd}$ in terms of $SL$, assuming the average context length per request is $\overline{N_{c}}$ and batch size is $BS$.
Lemma 4.1.
The total speculative decoding time $T_{sd}$ is a quadratic function of the speculative length $SL$:
(5) $\begin{split}T_{sd}&=BS\big(\frac{SL^{2}(\alpha_{d}+\alpha_{v})}{2}\\ &\quad+SL(\alpha_{d}\overline{N_{c}}+\alpha_{v}\overline{N_{c}}-\frac{\alpha_{d}}{2}+\frac{\alpha_{v}}{2}+\gamma_{d}+\gamma_{v}+\frac{\delta_{d}}{BS})\\ &\quad+\alpha_{v}\overline{N_{c}}+\gamma_{v}+\frac{\delta_{v}}{BS}\big)\end{split}$
Proof.
From Eq.  (1), the total speculative decoding time is $T_{sd}=T_{d}+T_{v}$. The expressions for $T_{d}$ and $T_{v}$ are given in Eqs.  (3) and  (4).
To express $T_{sd}$ in terms of speculative length $SL$ and batch size BSBS, we replace the summations $\sum_{i}N_{c}(i)$ and $\sum_{i}N_{b}(i)$ in Eqs.  (3) and  (4) using the average context length $\overline{N_{c}}$ and constant batch size. After simplification, we obtain the quadratic form in SLSL. ∎
This quadratic structure implies that decoding time grows super-linearly with speculative length, and thus, optimizing $SL$ is essential for balancing accuracy and latency.
@@
Figure 7. Average acceptance rates for different confidence score intervals of the draft model. The red dashed line connects (0,0) and (1,1) to aid in visual assessment.
4.2. The Number of Accepted Tokens
To quantify the effectiveness of speculative decoding, we develop a model to estimate the number of accepted tokens. Due to varying system loads and request types (see §2.3), we aim to provide a general estimation that avoids relying on request-specific conditions.
Let $BS$ denote the batch size and $SL$ the speculative length. The number of accepted tokens $NAT$ is defined as:
(6) $NAT=\sum_{i=1}^{BS}\left(\sum_{j=1}^{SL}AR_{i,j}+1\right)=\sum_{i=1}^{BS}\sum_{j=0}^{SL}AR_{i,j}$
@@ 4.2. The Number of Accepted Tokens
Here, $AR_{i,j}$ denotes the probability that the $j^{th}$ draft token of the $i^{th}$ request is accepted. To unify speculative and autoregressive decoding, we define $AR_{i,0}=1$, representing a bonus token that is always accepted.
Modeling acceptance rate. Acceptance in speculative decoding occurs when the draft model and the target model generate identical tokens at a given position, which implicitly implies high confidence from both models. Empirical observations (see Figure 7) confirm a strong correlation between confidence and acceptance. Similar to previous works (Du et al., 2024; Li et al., 2024), this finding suggests that we can model acceptance rates using confidence levels.
The probability that the $j^{th}$ draft token of the $i^{th}$ request is accepted depends on the successful acceptance of all previous tokens, and is approximated as:
(7) $AR_{i,j}=\prod_{k=0}^{j}p_{i,k}\approx\prod_{k=0}^{j}c_{i,k}$
where $p_{i,k}$ is the conditional probability of the $k^{th}$ draft token of the $i^{th}$ request is accepted given that all previous draft tokens have been accepted, and $c_{i,k}$ is its confidence score. Notably, we define $p_{i,0}=c_{i,0}=1$ to unify speculative and autoregressive decoding.
Figure 8. The relationship between estimated and actual number of accepted token with speculative length.
@@
The properties of the number of accepted tokens. We analyze the mathematical properties of the average number of accepted tokens $\overline{NAT}$, based on the confidence-based acceptance model.
Lemma 4.2.
The average number of accepted tokens $\overline{NAT}$ has the following properties:
  1. (1)
    (Initialization) When $SL=0$, $\overline{NAT}=1$;
  2. (2)
    (Marginal gain) For $SL\geq 1$, the marginal gain is $\Delta\overline{NAT}\in[0,1]$;
  3. (3)
@@
    (Diminishing returns) The marginal gain $\Delta\overline{NAT}$ is non-increasing as $SL$ increases.
Proof.
These results follow from the definition of NAT{NAT} in Eq. (6) and the confidence-based approximation of $AR_{i,j}$ in Eq. (7).
  • (1)
    When $SL=0$, only $AR_{i,0}=1$ contributes to the sum, hence $\overline{NAT}=1$.
  • (2)
    Increasing $SL$ by one adds an additional term $\Delta\overline{NAT}=\overline{AR_{SL}}=\frac{1}{BS}\sum_{i=1}^{BS}\prod_{k=0}^{SL}c_{i,k}\in[0,1]$, since each $c_{i,k}\in[0,1]$.
  • (3)
    For each $i$ and $j$, we have $AR_{i,j+1}=AR_{i,j}\cdot c_{i,j+1}\leq AR_{i,j}$ because $0\leq c_{i,j+1}\leq 1$. Therefore, $AR_{i,j}$ is non-increasing in $j$, and so is its average $\overline{AR_{j}}$. Thus, $\Delta\overline{NAT}=\overline{AR_{SL}}$ is non-increasing.
@@
4.3. The Relationship between Efficiency and Speculative Length
Let speculative length $SL$ be extended to a continuous variable $s\in\mathbb{R}_{\geq 0}$, and define the throughput function as:
(8) $f(s)=\frac{g(s)}{h(s)}$
@@ 4.3. The Relationship between Efficiency and Speculative Length
where $g(s)$ is the average number of accepted tokens per request (i.e., $\overline{NAT}$), and $h(s)$ is the average speculative decoding time per request (i.e., $T_{sd}/BS$).
Existence and location of optimal speculative length. Due to the complexity of $f(s)$, we analyze its properties indirectly by leveraging the known properties of $g(s)$ and $h(s)$.
Theorem 4.3.
$f(s)$ is quasi-concave on $\mathbb{R}_{\geq 0}$ and attains a unique global maximum $s^{*}\in\mathbb{R}_{\geq 0}$. Consequently, $f(s)$ is either monotonically decreasing, or strictly increasing on $0,s^{*})$ and strictly decreasing on $(s^{*},\infty)$.
Proof.
By Lemma 4.2, the numerator $g(s)$, which represents the average number of accepted tokens per request, is non-decreasing and concave, i.e., $g^{\prime}(s)\geq 0$ and $g^{\prime\prime}(s)\leq 0$. By Lemma 4.1, the denominator $h(s)=T_{sd}/BS$ is a quadratic function with positive coefficients, and hence strictly convex and increasing for $s\geq 0$.
The throughput function is defined as $f(s)=g(s)/h(s)$. Using standard quotient rule, its first derivative is:
(9) $f^{\prime}(s)=\frac{g^{\prime}(s)h(s)-g(s)h^{\prime}(s)}{\big(h(s)\big)^{2}}$
and the second derivative is:
(10) $\begin{split}f^{\prime\prime}(s)=&\frac{1}{\big(h(s)\big)^{4}}\Big(\big(g^{\prime\prime}(s)h(s)-g(s)h^{\prime\prime}(s)\big)\big(h(s)\big)^{2}\\ &-2h(s)h^{\prime}(s)\big(g^{\prime}(s)h(s)-g(s)h^{\prime}(s)\big)\Big)\end{split}$
Based on the derivative properties of $g(s)$ and $h(s)$, it follows that: (1) $f^{\prime}(s)$ changes sign at most once from positive to negative, implying that $f(s)$ is either strictly decreasing or unimodal (increasing then decreasing) with a single peak; (2) $f(s)$ is quasi-concave on $s\in\mathbb{R}_{\geq 0}$ and attains a unique global maximum $s^{*}\in\mathbb{R}_{\geq 0}$.. ∎
Specifically, the monotonicity of $f(s)$ is determined by the location of the optimal speculative length $s^{*}$.
Corollary 4.4.
The optimal speculative length $s^{*}$ is positively correlated with the acceptance and the computational gap between the target and draft models, and negatively correlated with the batch size $BS$ and the average context length $\overline{N_{c}}$.
Proof.
From the derivative expression in Theorem 4.3, we know that $f^{\prime}(0)>0$ is equivalent to $g^{\prime}(0)c-b>0$, which determines whether the optimal length $s^{*}>0$.
By substituting $c=h(0)$ and $b=h^{\prime}(0)$ from Lemma 4.1, we obtain:
@@
(11) $\begin{split}g^{\prime}(0)c-b=&2\alpha_{v}\overline{N_{c}}g^{\prime}(0)+2\gamma_{v}g^{\prime}(0)+\frac{2\delta_{v}}{BS}g^{\prime}(0)-2\alpha_{v}\overline{N_{c}}\\ &-\alpha_{v}-2\gamma_{v}-2\alpha_{d}\overline{N_{c}}+\alpha_{d}-2\gamma_{d}-\frac{2\delta_{d}}{BS}\end{split}$
From this expression, we observe: (1) A higher acceptance rate (which increases $g^{\prime}(0)$) contributes positively, as it scales all positive terms; (2) Reducing the batch size $BS$ increases the impact of both $\delta_{v}$ and $\delta_{d}$ terms due to their inverse dependence; (3) A shorter average context length $\overline{N_{c}}$ decreases the cumulative weight of the $\alpha_{v}$ and $\alpha_{d}$ coefficients; (4) Enlarging the computational gap (i.e., increasing $(\alpha_{v},\gamma_{v},\delta_{v})$ and decreasing $(\alpha_{d},\gamma_{d},\delta_{d})$) also favors a higher value of $g^{\prime}(0)c-b$.
Hence, the optimal $s^{*}$ shifts to the right under these conditions. ∎
Remark. The structural property of $f(s)$ in Theorem 4.3 enables a practical tuning method. Since $f(s)$ is either monotonically decreasing or unimodal with a unique maximum, the optimal speculative length $s^{*}$ can be located by incrementally evaluating $f(s)$ at increasing speculative length and stopping at the first local maximum. This allows lightweight online tuning without exhaustive search.
5. Adaptive Speculative Decoding